A sample of 137 burglaries indicated that 58 were not via an open or unlocked door window. At the 0.05 level of significance, test the claim that less than 50% of the burglaries are not through an open or unlocked door or window.
Answer #1

Hypotheses:

Ho: p = 0.50

Ha: p < 0.50 [Left Tailed test]

Sample proportion = 58/137 = 0.423

Test statistic

z = -1.80

p - Value = P(z < -1.80) = 0.036

Since p - value is less than 0.05, we reject Ho. Hence,

There is sufficient evidence to support the claim that less than 50% of the burglaries are not through an open or unlocked door or window.